1. Boeing is expected to pay an annual dividend of $3.75 this coming year. The stock is selling for $100.90 a share and has a required return of 11 percent. What is the growth rate of the dividend?

a) 4.72 percent

b) 7.2 percent

c) 3.72 percent

d) 3.62 percent

e) 14.72 percent

2. The common stock of BBY is valued at $28.76 a share. The company increases its dividend by 6.5 percent annually and expects its next dividend to be $0.84 per share. What is the required rate of return on this stock?

a) 6.64 percent

b) 7.53 percent

c) 9.42 percent

d) 10.17 percent

e) 8.92 percent
Answer #1

Required return=(D1/Current price)+Growth rate

1.

0.11=(3.75/100.90)+Growth rate

Growth rate=0.11-(3.75/100.90)

=7.2%(Approx).

2.Required return=(0.84/28.76)+0.065

=9.42%(Approx).